Find first-order partial derivatives of F(x)= cosx²y²z²/Arc sinxy²


1
Expert's answer
2020-04-14T17:37:34-0400

Consider the function "f(x,y,z)=\frac{\cos(x^2y^2z^2)}{\arcsin(xy^2)}"


Differentiate the function partially with respect to "x" as,



"f_x(x,y,z)=\frac{\partial}{\partial x}(\frac{\cos(x^2y^2z^2)}{\arcsin(xy^2)})"


"=\frac{1}{\arcsin^2(xy^2)}(\arcsin(xy^2)\frac{\partial}{\partial x}(\cos(x^2y^2z^2))-(\cos(x^2y^2z^2))\frac{\partial}{\partial x}(\arcsin(xy^2)))"


"=\frac{-2xy^2z^2\sin(x^2y^2z^2)\arcsin(xy^2)-\frac{y^2\cos(x^2y^2z^2)}{\sqrt{1-x^2y^4}}}{\arcsin^2(xy^2)}"



"=\frac{-2xy^2z^2\sin(x^2y^2z^2)\arcsin(xy^2)\sqrt{1-x^2y^4}-y^2\cos(x^2y^2z^2)}{\arcsin^2(xy^2)\sqrt{1-x^2y^4}}"

Differentiate the function partially with respect to "y" as,



"f_y(x,y,z)=\frac{\partial}{\partial y}(\frac{\cos(x^2y^2z^2)}{\arcsin(xy^2)})"


"=\frac{1}{\arcsin^2(xy^2)}(\arcsin(xy^2)\frac{\partial}{\partial y}(\cos(x^2y^2z^2))-(\cos(x^2y^2z^2))\frac{\partial}{\partial y}(\arcsin(xy^2)))"


"=\frac{-2x^2yz^2\sin(x^2y^2z^2)\arcsin(xy^2)-\frac{2xy\cos(x^2y^2z^2)}{\sqrt{1-x^2y^4}}}{\arcsin^2(xy^2)}"



"=\frac{-2x^2yz^2\sin(x^2y^2z^2)\arcsin(xy^2)\sqrt{1-x^2y^4}-2xy\cos(x^2y^2z^2)}{\arcsin^2(xy^2)\sqrt{1-x^2y^4}}"



Differentiate the function with respect to "z" as,



"f_z(x,y,z)=\frac{\partial}{\partial z}(\frac{\cos(x^2y^2z^2)}{\arcsin(xy^2)})"



"=\frac{1}{\arcsin(xy^2)}\frac{\partial}{\partial z}(\cos(x^2y^2z^2))"



"=\frac{-2x^2y^2z\sin(x^2y^2z^2)}{\arcsin(xy^2)}"

Therefore, the partial derivatives of the multi variable function are:

"f_x(x,y,z)=\frac{-2xy^2z^2\sin(x^2y^2z^2)\arcsin(xy^2)\sqrt{1-x^2y^4}-y^2\cos(x^2y^2z^2)}{\arcsin^2(xy^2)\sqrt{1-x^2y^4}}"

"f_y(x,y,z=\frac{-2x^2yz^2\sin(x^2y^2z^2)\arcsin(xy^2)\sqrt{1-x^2y^4}-2xy\cos(x^2y^2z^2)}{\arcsin^2(xy^2)\sqrt{1-x^2y^4}}"


"f_z(x,y,z)=\frac{-2x^2y^2z\sin(x^2y^2z^2)}{\arcsin(xy^2)}"